OpenStudy (anonymous):

For three events A, B and C , P(exactly one of the event A or B occurs) = P(exactly one of the event B or c occurs) = P(exactly one of the event C or A occurs) = p and P(all the three avents occur simultaneously) = P^2 , where 0

OpenStudy (kropot72):

P(A) = P(B) = P(C) P(A + B + C) = P^2 \[P ^{2}=P(A)\times P(B)\times P(C)=P(A)^{3}\] \[P(A)=\sqrt[3]{P ^{2}}\]